The fission of uranium nuclei in a nuclear reactor produces high speed neutrons. Before a neutron can trigger additional fissions, it has to be slowed down by collisions with carbon nuclei in the moderator of the reactor. Suppose a neutron with mass of 1.0u travelling at 2.0x107m/s undergoes a head-on elastic collision with a carbon nucleus with mass of 12u initially at rest. Calculate the velocities of the neutron and carbon after the collision.



1
Expert's answer
2021-06-11T11:21:36-0400

Let us write the conservation of momentum law:

"m_n v_{n,0} + m_cv_{c,0} = m_n v_{n,1} + m_cv_{c,1}" where "v_{c,0}=0."

"m_n v_{n,0} = m_n v_{n,1} + m_cv_{c,1}"

Due to elastic collision, the energy is conserved, so

"\dfrac{m_n v_{n,0}^2}{2} + \dfrac{m_cv_{c,0} ^2}{2}= \dfrac{m_n v_{n,1}^2}{2} + \dfrac{m_cv_{c,1}^2}{2}" or "m_n v_{n,0}^2 = m_n v_{n,1}^2 + m_cv_{c,1}^2".

So we have a system of equations

"\begin{cases}
m_n v_{n,0} = m_n v_{n,1} + m_cv_{c,1},\\
m_n v_{n,0}^2 = m_n v_{n,1}^2 + m_cv_{c,1}^2
\end{cases},"

"\begin{cases}
1 v_{n,0} = 1 v_{n,1} + 12v_{c,1},\\
1 v_{n,0}^2 = 1 v_{n,1}^2 + 12v_{c,1}^2
\end{cases}"

We solve these equation by substituting "v_{n,1} =v_{n,0} - 12v_{c,1}" into the second equation and get "v_{c,1} = \dfrac{2}{13}v_0 = \dfrac{2}{13}\cdot 2.0\cdot10^7\,\mathrm{m/s}= 3.1\cdot10^6\,\mathrm{m/s}" .

"v_{n,1} = v_{n,0} - 12\cdot v_{c,1} = -1.7\cdot10^7\,\mathrm{m/s}."